OpenStudy (logo):

Write a recursive formula for the sequence 5, 18, 31, 44, 57, … Then find the next term. help pls! Will award medal!!!

OpenStudy (royalranger):

\[t_n = t_{n-1} +13\]

OpenStudy (sunnnystrong):

If the sequence is 5, 18, 31, 44, 57, … Than we are adding some number to each step. \[a _{n}=a_{n-1}+d\] an= nth term in sequence an-1=term before the nth term d= common difference

OpenStudy (sunnnystrong):

^^^ So what @royalranger said would be correct :D

OpenStudy (royalranger):

@sunnnystrong equation works just as well. I think you can find the last term.

OpenStudy (sunnnystrong):

In function notation: t(n)=t(n-1)+13

OpenStudy (logo):

ah I get it.. So then t(n) = 5+ 13 (n - 1)

OpenStudy (logo):

the next term is 70 right?

OpenStudy (royalranger):

Yes it is 70.

OpenStudy (sunnnystrong):

yep!

OpenStudy (sunnnystrong):

good job @logo :)

OpenStudy (logo):

@royalranger @sunnnystrong thx!

OpenStudy (royalranger):

No problem.